- Status Report
- August 8, 2022
SATOPS Morning Report 08-02-2000
SATOPS MORNING REPORT – August 2, 2000
No change in the status of the GOES spacecraft.
GOES-8 rapid scan operations were activated from 1726Z to 0026Z for
severe weather over Wisconsin.
GOES-10 frame breaks occurred at 0310Z and 0536Z due to loss of signal
synchronization, and at 0334Z and 0406Z due to thunderstorms at the Wallops
No change in the status of the POES spacecraft.
NOAA-15 HIRS continues to display erratic behavior. Filter wheel
motor current is still within prescribed limit guidelines.
NOAA-12 rev 47871, G4A was short due to Fairbanks CDA configuration.
100% of the data was recovered.
NOAA-14 rev 28803, G5B noisy from the spacecraft with 100% of the data
NOAA-12 rev 47874, L4A did not lock on bit synch. 100% of the data was
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 24 hours.
Questions or comments to email@example.com